# A Constant Measurement Quantum Algorithm for Graph Connectivity

### v1

@@ III.2 ZX-calculus
The ZX-calculus is a graphical language to describe quantum circuits and, more generally, linear maps [35, 36]. For an introduction coming from a quantum computing background, we refer to [37], for the perspective from its category theory roots, we refer to [38]. A ZX-diagram is made of spiders representing multi-dimensional tensors in either $Z$ or $X$ basis which are connected via wires. They are defined as:
$:=|0\rangle^{\otimes m}\langle 0|^{\otimes n}+e^{i\alpha}|1\rangle^{\otimes m}\langle 1|^{\otimes n}$ (1a)
$:=|-\rangle^{\otimes m}\langle-|^{\otimes n}+e^{i\alpha}|+\rangle^{\otimes m}\langle+|^{\otimes n}$ (1b)
@@ III.3 Non-unitary gates
where $N$ is a non-unitary transformation, in the present case the white spider of equation (1). The probability of success upon measurement is given as [41]:
@@
with a normalization constant $c$. The success probability relates to the eigenvalues of $N$. From definition of (1a), it is apparent that it only has eigenvalues of 00 and $1$. For a given state $|\psi\rangle$, the Z spider has success probabilities
@@
Larger spiders follow the same pattern, with a matrix constructed with $1$ in the first entry, $\exp(i\alpha)$ in the last one.
[…]
@@ IV.1 Construction
The contraction of the individual two-qubit spiders can be understood iteratively through the two possible cases. Either two spiders are disconnected and they cannot be contracted.
[…]
When they do share at least one wire, they can be contracted together to a single spider. It does not matter how many wires connect to the spider, leading to successively larger spiders equivalent to the network of initial two-qubit spiders.
[…]
@@ IV.4 State decay
up to the phase $\alpha$. Whereas stabilizers due to their unitary nature keep all other states alive, at most transforming them, the non-unitary gate dissipates their information to the open environment and that part of the state is lost.
@@
[…]
@@ IV.6 Experimental verification
To best of our knowledge, there is no quantum computer that implements the required non-unitary operation. However, there are a number of experiments that create large GHZ states. Experimentally, a the creation of GHZ state sizes has been pushed in recent years [48, 49, 50] and experiments with 1000’s of qubits have been proposed [51]. In the case of applications on commercial hardware, the GHZ state is generally created using CNOT and CZ operations. The large scale of the experiments show that the result of our algorithm is generally achievable even on near-term hardware.
@@
[…]
where the box between the spiders is the Hadamard gate. This difference prevents the contraction to a single object.
@@ Appendix A Simulation results
As the necessary operations are not currently implemented in a quantum computer, we show the structure of the algorithm in simulation. Any state simulation-based program should be able to perform the necessary calculation. We implement our system in Julia [56] and its quantum simulation package Yao [57]. In the supplementary material, we provide a notebook that implements an interface between a graph and the quantum algorithm. The circuit is implemented without ancilla qubits and ignores state decay.
@@ B.1 Upper bound as an integer partition problem
Table 3: The structure of possible partitions $\lambda$ within a row of length $n^{\prime}$. The last line indicates the some over the row.
@@
For example, when $n=5$ and $m=10$, the largest possible depth is $d=10$. In this case, the only possible partition of $m=10$ is $\lambda=(1,1,1,1,1,1,1,1,1,1)$, which corresponds the following Young diagram,
[Uncaptioned image] (35)
Then one obtains $N_{10}=10\times\binom{5}{2}=100$. For $d=9$, the only possible partition of $m=10$ is $\lambda=(2,1,1,1,1,1,1,1,1)$, and thus the corresponding Young diagram is
[Uncaptioned image] (36)
Then one obtains $N_{9}=9\times(\binom{5}{4}+8\times\binom{5}{2})=765$. For $d=8$, the two possible partition of $m=10$ is $\lambda=(3,1,1,1,1,1,1,1)$ and $(2,2,1,1,1,1,1,1)$, and thus the corresponding Young diagrams are
[Uncaptioned image] (37)
Due to the restriction $2\lambda_{i}\leq n$ for all $\lambda_{i}$, the largest possible value of $\lambda_{i}$ should be $\left\lfloor\frac{n}{2}\right\rfloor$. In our case, the largest $\lambda_{i}$ is $2$. Hence, the first Young diagram does not contribute. Counting the second Young diagram, we obtain $N_{8}=\binom{8}{2}\times(2\times\binom{5}{4}+6\times\binom{5}{2})=1960$.
Using the same logic, for $d=7$, although there are 3 partitions, the only relevant partition for $m=10$ is $\lambda=(2,2,2,1,1,1,1)$, and its corresponding Young diagram is
[Uncaptioned image] (38)
Then one obtains $N_{7}=\binom{7}{3}\times(3\times\binom{5}{4}+4\times\binom{5}{2})=1925$. For $d=6$, while there are a total 5 partitions, the only relevant partition for $m=10$ is $\lambda=(2,2,2,2,1,1)$, and its corresponding Young diagram is
[Uncaptioned image] (39)
Then one obtains $N_{6}=\binom{6}{4}\times(4\times\binom{5}{4}+2\times\binom{5}{2})=600$. Finally, for $d=5$, while there are a total of 7 partitions, the only relevant partition for $m=10$ is $\lambda=(2,2,2,2,2)$, and its corresponding Yong diagram is
[Uncaptioned image] (40)

### v2

@@ III.2 ZX-calculus
The ZX-calculus is a graphical language to describe quantum circuits and, more generally, linear maps [35, 36]. For an introduction coming from a quantum computing background, we refer to [37], for the perspective from its category theory roots, we refer to [38]. A ZX-diagram is made of spiders representing multi-dimensional tensors in either $Z$ or $X$ basis which are connected via wires. The definition is given in figure 1.
[…]
Refer to caption$:=|-\rangle^{\otimes m}\langle-|^{\otimes n}+e^{i\alpha}|+\rangle^{\otimes m}\langle+|^{\otimes n}$
Figure 1: The main elements of ZX calculus are the spiders. Contrary to standard quantum circuit elements, the number of inputs and outputs does not have to be fixed.
@@ III.3 Non-unitary gates
where $N$ is a non-unitary transformation, in the present case the white spider of figure 1. The probability of success upon measurement is given as [41]:
@@
with a normalization constant $c$. The success probability relates to the eigenvalues of $N$. From definition of a spider given in figure 1, it is apparent that it acts as a projector. For a given state $|\psi\rangle$, the Z spider has success probabilities
@@
$Z(\alpha)=\begin{pmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&e^{i\alpha}\end{pmatrix}$ (6)
Larger spiders follow the same pattern, with a matrix constructed with $1$ in the first entry, $\exp(i\alpha)$ in the last one, as shown in figure 4.
@@
Figure 4: The general matrix representation of a multi-legged Z spider. Only the first and last entry of the matrix have an entry, $1$ and $exp(i\alpha$ respectively.
@@ IV.1 Construction
The contraction of the individual two-qubit spiders can be understood iteratively through the two possible cases. Either two spiders are disconnected and they cannot be contracted. When they do share at least one wire, they can be contracted together to a single spider. It does not matter how many wires connect to the spider, leading to successively larger spiders equivalent to the network of initial two-qubit spiders.
@@ IV.4 State decay
up to the phase $\alpha$. Whereas stabilizers due to their unitary nature keep all other states alive, at most transforming them, the non-unitary gate dissipates their information to the open environment and that part of the state is lost. Ignoring this information loss can lead to nonsensical results. It is possible to construct quantum algorithms that solve NP problems with a polynomial runtime, if the state decay is ignored [48].
@@
Figure 6: A possible translation from the non-unitary spider to unitary operations together with a measurement. The $M_{X}$ box indicates a measurementi in the X basis, such that the state on the two continuing qubits is a Bell state, $|00\rangle\pm|11\rangle$.
This is shown in figure 6. The measurement in the X basis projects the two other qubits of the GHZ state to a Bell state $|00\rangle\pm|11\rangle$, without loss of information to the outside environment. Measuring in the Z basis removes the entanglement [1].
@@ IV.6 Experimental verification
To best of our knowledge, there is no quantum computer that implements the required non-unitary operation. However, there are a number of experiments that create large GHZ states. Experimentally, a the creation of GHZ state sizes has been pushed in recent years [49, 50, 51] and experiments with 1000’s of qubits have been proposed [52]. In the case of applications on commercial quantum computers, the GHZ state is generally created using CNOT and CZ operations. The large scale of the experiments show that the result of our algorithm is generally achievable even on near-term hardware.
@@ Appendix A Simulation results
As the necessary operations are not currently implemented in a quantum computer, we show the structure of the algorithm in simulation. Any state simulation-based program should be able to perform the necessary calculation. We implement our system in Julia [57] and its quantum simulation package Yao [58]. In the supplementary material, we provide a notebook that implements an interface between a graph and the quantum algorithm. The circuit is implemented without ancilla qubits and ignores state decay, as measurement samples are taken from the remaining state. The decay can be observed when inspecting the state.
@@ B.1 Upper bound as an integer partition problem
Table 3: The structure of possible partitions $\lambda$ within a row of length $n^{\prime}$. The last line indicates the sum over the row.
@@
Figure 7: The Young diagrams used to calculate the example of $n=5$ and $m=10$.
For example, when $n=5$ and $m=10$, the largest possible depth is $d=10$. In this case, the only possible partition of $m=10$ is $\lambda=(1,1,1,1,1,1,1,1,1,1)$. Then one obtains $N_{10}=10\times\binom{5}{2}=100$. For $d=9$, the only possible partition of $m=10$ is $\lambda=(2,1,1,1,1,1,1,1,1)$ Then one obtains $N_{9}=9\times(\binom{5}{4}+8\times\binom{5}{2})=765$.
For $d=8$, the two possible partition of $m=10$ is $\lambda=(3,1,1,1,1,1,1,1)$ and $(2,2,1,1,1,1,1,1)$. Due to the restriction $2\lambda_{i}\leq n$ for all $\lambda_{i}$, the largest possible value of $\lambda_{i}$ should be $\left\lfloor\frac{n}{2}\right\rfloor$. In our case, the largest $\lambda_{i}$ is $2$. Hence, the first Young diagram does not contribute. Counting the second Young diagram, we obtain $N_{8}=\binom{8}{2}\times(2\times\binom{5}{4}+6\times\binom{5}{2})=1960$.
Using the same logic, for $d=7$, although there are 3 partitions, the only relevant partition for $m=10$ is $\lambda=(2,2,2,1,1,1,1)$. Then one obtains $N_{7}=\binom{7}{3}\times(3\times\binom{5}{4}+4\times\binom{5}{2})=1925$. For $d=6$, while there are a total 5 partitions, the only relevant partition for $m=10$ is $\lambda=(2,2,2,2,1,1)$.
Then one obtains $N_{6}=\binom{6}{4}\times(4\times\binom{5}{4}+2\times\binom{5}{2})=600$. Finally, for $d=5$, while there are a total of 7 partitions, the only relevant partition for $m=10$ is $\lambda=(2,2,2,2,2)$. All Young diagrams mentioned here are drawn in figure 7.